OpenStudy (anonymous):

Solve dy\dx + xsin(2y) = x^3cos^2y

OpenStudy (anonymous):

\[\begin{align*} \frac{dy}{dx}+x\sin2y&=x^3\cos^2y\\ (x\sin2y-x^3\cos^2y)~dx+dy&=0 \end{align*}\] Uh... any takers?

OpenStudy (anonymous):

\[M_y=2x\cos2y+2x^3\cos y\sin y=2x\cos2y+x^3\sin2y\\ N_x=0\] Hmm...

OpenStudy (anonymous):

Integrating factor(s): \[\mu(x)=\exp\left(\int\frac{M_y-N_x}{N}~dx\right)\\ \mu(x)=\exp\left(\int (2x\cos2y+x^3\sin2y)~dx\right)\] \[\eta(y)=\exp\left(\int\frac{N_x-M_y}{M}~dy\right)\\ \eta(y)=\exp\left(\int\frac{2x\cos2y+x^3\sin2y}{x^3\cos^2y-x\sin2y}~dy\right)\] \(\mu(x)\) would be easier to simplify but it's not in terms of \(x\) only... The equation definitely isn't separable, since \(M_y\not=N_x\), which I showed in @ameramd's previous question is a required condition for separability.

OpenStudy (anonymous):

My next guess would be some clever substitution, but not sure what it would be

OpenStudy (anonymous):

thanx @SithsAndGiggles it's bernoulli Eqn.

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (goformit100):

Giving you an energetic and warm Welcome to one and only 'Open Study'.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }+x \sin 2 y=x^3 \cos ^2y\] divide by \[\cos ^2y\] \[\sec ^2y \frac{ dy }{ dx }+x \frac{ 2\sin y \cos y }{ \cos ^2y }=x^3\] \[\sec ^2y \frac{ dy }{ dx }+2x \tan y=x^3\] put \[\tan y=t,\sec ^2y~\frac{ dy }{ dx }=\frac{ dt }{ dx }\] \[\frac{ dt }{ dx }+2 x t=x^3\] \[I.F=e ^{\int\limits 2x dx}=e ^{x^2}\] solution is \[t e ^{x^2}=\int\limits x^3e ^{x^2}dx+c\] \[t e ^{x^2}=\int\limits x^2 e ^{x^2}x dx+c=I1+c\] put\[x^2=z,2xdx=dz,x dx=\frac{ dz }{ 2 }\] \[I1=\frac{ 1 }{ 2 } \int\limits z e^z dz=\frac{ 1 }{ 2 }\left[ z e^z-\int\limits 1*e^z dz \right]=\frac{ 1 }{ 2}\left[ z e^z-e^z \right]\] \[=\frac{ 1 }{ 2 }\left( z-1 \right)e^z=\frac{ 1 }{ 2 }\left( x^2-1 \right)e ^{x^2}\] \[t e ^{x^2}=\frac{ 1 }{ 2 }\left( x^2-1 \right)e ^{x^2}+c\] \[\tan y~ e ^{x^2}=\frac{ 1 }{ 2 }\left( x^2-1 \right)e ^{x^2}+c\]

OpenStudy (anonymous):

y'sec^2y+2xtany=x3 u=tany u'+2xu=x3 u=exp(-x2)[c+int(x3exp(x2)dx]=exp(-x2)[c+x2/2exp(x2)-1/2exp(x2)] then tany=cexp(-x2)+1/2(x2-1)